Work through the simplex method (in algebraic form) step by step to solve the following problem.

Maximize Z = 4?1 + 3?2 + 6?3,
Subject to
3?1 + ?2 + 3?3 ≤ 30
2?1 + 2?2 + 3?3 ≤ 40
and
?1 ≥ 0, ?2≥ 0, ?3 ≥ 0.
Answer #1

Solution:
Problem is

Max Z = 4 x1 + 3 x2 + 6 x3
subject to
3 x1 + x2 + 2 x3 30
2 x1 + 2 x2 + 3 x3 40
and x1,x2,x3≥0;



The problem is converted to canonical form by adding slack, surplus and artificial variables as appropiate

1. As the constraint-1 is of type '≤' we should add slack variable S1

2. As the constraint-2 is of type '≤' we should add slack variable S2

After introducing slack variables

Max Z = 4 x1 + 3 x2 + 6 x3 + 0 S1 + 0 S2
subject to
3 x1 + x2 + 2 x3 + S1 = 30
2 x1 + 2 x2 + 3 x3 + S2 = 40
and x1,x2,x3,S1,S2≥0


Iteration-1 Cj 4 3 6 0 0
B CB XB x1 x2 x3 S1 S2 MinRatio
XB/x3
S1 0 30 3 1 2 1 0 30/2=15
S2 0 40 2 2 (3) 0 1 40/3=13.3333
Z=0 Zj 0 0 0 0 0
Zj-Cj -4 -3 -6↑ 0 0



Negative minimum Zj-Cj is -6 and its column index is 3. So, the entering variable is x3.

Minimum ratio is 13.3333 and its row index is 2. So, the leaving basis variable is S2.

∴ The pivot element is 3.

Entering =x3, Departing =S2, Key Element =3

R2(old) = 40 2 2 3 0 1
R2(new)=R2(old)÷3 13.3333 0.6667 0.6667 1 0 0.3333



R1(new)=R1(old) - 2R2(new)

R1(old) = 30 3 1 2 1 0
R2(new) = 13.3333 0.6667 0.6667 1 0 0.3333
2×R2(new) = 26.6667 1.3333 1.3333 2 0 0.6667
R1(new)=R1(old) - 2R2(new) 3.3333 1.6667 -0.3333 0 1 -0.6667


Iteration-2 Cj 4 3 6 0 0
B CB XB x1 x2 x3 S1 S2 MinRatio
S1 0 3.3333 1.6667 -0.3333 0 1 -0.6667
x3 6 13.3333 0.6667 0.6667 1 0 0.3333
Z=80 Zj 4 4 6 0 2
Zj-Cj 0 1 0 0 2



Since all Zj-Cj≥0

Hence, optimal solution is arrived with value of variables as :
x1=0,x2=0,x3=13.3333

Max Z=80

Please comment if any doubt

Rate me

Thanks